OpenStudy (megannicole51):

Find the first 4 terms of the taylor series for the function about the point a. 1/sqrt(1+x)

OpenStudy (dumbcow):

this means we need first 3 derivatives of the function....they dont give a value for "a" ?

OpenStudy (anonymous):

First 4 derivatives.

OpenStudy (dumbcow):

@Lessis , no 1st term is just original function

OpenStudy (megannicole51):

im sorry its not a its 0....ive been doing this stupid math for way too long:/

OpenStudy (anonymous):

Usually, the first term of the taylor series isn't counted, since it's just f(a) (And the taylor polinomial of grade n is f(a) plus the first n derivatives). Derivating that is going to be a chore though.

OpenStudy (dumbcow):

@megannicole51 haha ok thought so

OpenStudy (megannicole51):

its seriously been a long day/night my bad!

OpenStudy (megannicole51):

lol! nice catch @dumbcow

OpenStudy (dumbcow):

ok what is 1st derivative?

OpenStudy (megannicole51):

-1/2(x+1)^(3/2)

OpenStudy (megannicole51):

second derivative is ((3)/(4(x+1)^(5/2)))

OpenStudy (dumbcow):

yes...exponents are neg right

OpenStudy (megannicole51):

no

OpenStudy (dumbcow):

should be , you started with power of (-1/2) then each time you take derivative you subtract 1

OpenStudy (megannicole51):

oh ok

OpenStudy (anonymous):

It really depends on how you're writing it. If your power is already in the denominator, there is no need to add a minus sign to it.

OpenStudy (megannicole51):

okay

OpenStudy (megannicole51):

now what?

OpenStudy (dumbcow):

sub in 0 for x then use formula for taylor series you still need 3rd derivative i think

OpenStudy (megannicole51):

f'(0)=-1/2 f''(0)=3/4 f'''(0)=-15/8 f''''(0)=105/16

OpenStudy (dumbcow):

yep

OpenStudy (dumbcow):

Taylor series \[1-\frac{1}{2}x+\frac{3}{8}x^{2}-\frac{5}{16}x^{3}\]

OpenStudy (megannicole51):

yay:) thank you!

OpenStudy (dumbcow):

yw